a solution is obtained by mixing 300g of 25% solution and 400g of 40% solution by mass .calculate the mass% of the resulting mixture

Grade:10

2 Answers

Arun
5 years ago
Dear student

According to question,
300g of 25% solution contains solute = (300 x 25) / 100 = 75g
&
400g of 40% solution contains solute = (400 x 40) / 100 = 160g
Total solute = 75 + 160 = 235g
As given in the question, total solution =300 + 400 = 700g
Percentage of solute in the final solution = (235 x100) / 700 = 33.5%
Percentage of water in the final solution = 100 - 33.5 = 65.5%
prince sunjot dutt
5 years ago
Mass% =(mass of one component/total mass of solution)×100
Acc. Ques:
In 25% solution, mass of solute=( 25×300)\100=75 g
Mass of solvent=300-75=225g
Similarly,
In 40% solution, Mass of solute= 40×400)100= 160g 
Mass of solvent=240g
Therefore, total solute=235g
Total solvent=465g
Mass%=(235/465)*100=50.53%